OpenStudy (anonymous):

y=ae^(bx) for points (3,10) and (0,1)

OpenStudy (anonymous):

here substituting for (0,1) gives a=1 again substituting for (3,10) gives 10=e^3b hence b=(1/3 )(ln10)

OpenStudy (anonymous):

so y= 10e^(3b)?